Home‎ > ‎Reports‎ > ‎crew114‎ > ‎

day02


Crew #114 Log Book for February 26, 2012

Photos of the day





Commander's Check-In Report

Date and time: February 26, 2012, 19:00

Written by: Balwant Rai, Commander, Crew 114

Crew Physical Status: Excellent

Time Departed/Returned from EVA: see EVA reports

Brief Narrative of Field Mission Results: All the medical, geological, rover, and Human factors experiments started today with interesting results.

EVA Data/Interpretations:We completed two successful EVA’s. Many lessons were learned, that included: defogging of helmet before going for EVA and proper use of the EVA suit backpacks

Engineering/Hab Maintenance: Need more generator fuel, Internet too slow

Report Transmission Schedule: by 20:00

Plans for Tomorrow: Will begin with doing the medical, human factors, geology and rover experiments again. Soil samples will again be collected. Medical data will be collected during two EVA’s; one will be on foot, and in other the ATV’s will be used. Tomorrow will be a non cooking day.

Inventory:We ran out of baking soda.

Support Requested: Need more baking soda.

Miscellaneous: We were very creative, since our rover engineer, emergency physician, HSO and crew geologist began to really get their experiments started. The meals today were especially delicious, since it was a cooking day on Mars.


Commander's Report


Date and time: February 26, 2012, 19:00

Written by: Balwant Rai, Crew 114, Commander

Today we started with a delicious breakfast of cereals, ice tea and coffee planned by Health and safety officer and Chief engineer. It was very enjoyable and delicious. Afterwards, we began medical and human factors experiments.

Today was a great day for the crew to explore the Martian landscape and practice using the NASA rover. Afterwards, we had a great lunch prepared by Jasdeep kaur and Pierrick Darneaux. This was followed by two EVA’s that involved collecting soil samples for analysis of microbiological activity and to explore the Martian landscape.

During the late afternoon, the cadets worked on getting the rover working and examining how well it maneuvers when being controlled on the outside of the Hab. Our chefs (Jasdeep and Pierrick Darneaux) are currently working on a good cooked meal tonight, since it is a cooking-day. 


Tomorrow, we plan to begin with doing the medical and human factors experiment again on all crew members. The plan is to do this every day for the entire time we are here on Mars. The rover engineer, Yannick Wable, is planning on continuing his study of the rover and potentially using the drone. Pierrick and Thomas Quemin will also join Pierrick in examining the feasibility of the rover here on Mars. Wouter Poos will continue his collection of soil samples. Medical data will be collected during two EVA’s tomorrow. One will be on foot, and one will use the ATV’s. Tomorrow will be a non cooking day, so we look forward for cold meals.


Science Report (Geology) 


Date and time: February 26, 2012 11:00

Written by: Wouter Poos, Crew 114, Crew Geologist

Took surface and superficial microbiological samples at site 1 during EVA 1. No problems encountered. Samples are stored in the refrigerator for further analysis.

Science Report (Rover) 

Date: February 26, 2012

Written by: Wable Yannick, Rover Engineer, Crew 114

The executive officer (Quemin Thomas), chief engineer (Darneaux Pierrick), and the rover engineer (Wable Yannick) completed the first stage to understand how to drive the Rover. We successfully achieved our goal by driving the Rover with and without indications.

Engineering Report

Date and time: 26 February, 2012, 17:30
Written by Pierrick DARNEAUX, Crew 114, Crew Engineer.

Generator/Electricity:
Did you have any electricity outage? (Yes/No): NO

Fuel Status:
Diesel (Full/0.75/0.5/0.25/Empty): <0.25
Propane Tank (Full/0.75/0.5/0.25/Empty): <0.25
Gasoline Tank (Full/0.75/0.5/0.25/Empty): <0.25

Oil:
ATV Oil Quantity (# of Quarts in storage): 20
Car Oil Quantity (# of Quarts in storage): N/A

Water Status:
Outside Potable Water Tank Level (inches from bottom): 12
Trailer Potable Water Tank Level (inches from bottom): 14
Hab Potable Water Tank Level (inches from bottom): 7
Potable Water Meter Reading: 0051534.4
Into GreenHab Meter Reading: N/A
Notes/Comments:

GreenHab:
Condition of plants in Tank 1: N/A
Condition of plants in Tank 2: N/A
Condition of plants in Tank 3: N/A

Observatory:
Telescope Used? (Yes/No): YES
Observatory Notes/Comments: N/A

Transportation:
Hab Car (NOT RENTAL)
Used (yes/no): N/A
Oil Checked (yes/no): YES: N/A
Tire Status: N/A
Odometer at end of day: N/A
Notes/Comments on Hab Car: No car

ATV 1 (Opportunity):
Used: NO
Oil Checked: YES: Full
Fuel Consumed: 0
Tire Status: GOOD
Today's Run Time: 0
Season's Run Time: 87:55
Comments:-

ATV 2 (Spirit):
Used: NO
Oil Checked: YES: Full
Fuel Consumed: 0
Tire Status: GOOD
Today's Run Time: 0
Season's Run Time: 82:55
Comments:-

ATV 3 (Viking I):
Used: NO
Oil Checked: YES: Full
Fuel Consumed: 0
Tire Status: GOOD
Today's Run Time: 0
Season's Run Time: 85:55
Comments:-

Heating and Ventilation:
Thermostat setting upstairs (Farenheit):66°F
Actual temperature upstairs (Farenheit):64°F

Thermostat setting downstairs (Farenheit):65°F
Actual temperature downstairs (Farenheit):61°F

Computers/Networking Infrastructure:
Notes/comments on internet/computers?: still slow and low bandwidth

Miscellaneous: Electricity generating generators has a very low fuel level, please advise

General Engineering Notes/Comments: N/A


EVA 1 Report


Written by: Thomas Quemin, Crew 114, Executive Officer

Date and time: February 26, 2012 -19:00

Crew Members: Geologist (Woulter Poos who leads the experiment), Health and Safety Officer (Jasdeep Kaur), Executive Officer (Thomas Quemin)

Time: 10:00-11:20

Site Location: UTM Undisclosed

Transit mode: ATV

Objective: Take soil samples for analysis

Results: Samples taken at the EVA site

Lessons learned: Remember to defog helmet before going on EVA.

EVA 2 Report

Date and time: February 26, 2012, 19:00

Written by: Wouter Poos, Crew 114, Crew Geologist

Time: 15:30-16:30

Crew Members: Balwant Rai, Wouter Poos

Site Location: UTM undisclosed

Transit Mode: On Foot

Objectives: reconnaissance

Results: successful

Lessons Learned: Treating helmet with saliva reduces fogging problem, terrain north of the habitat is of sufficient cohesion for easy traverse.